# A central limit theorem for the stochastic cable equation

### v1

@@ Abstract
In this paper, we consider one-dimensional nonlinear stochastic cable equations driven by a multiplicative space-time white noise. Using the Malliavin–Stein’s method, we prove the Gaussian fluctuation for the spatial average of the solution under the total variation distance with mild conditions. We also establish functional central limit theorem.
@@ 1 Introduction and main results
We consider the nonlinear stochastic cable equation
@@
on $[0,T]\times[0,L]$ for some constants $\alpha\in\mathbb{R}$ and $\beta>0$, where $T>0$ is fixed, $\dot{W}$ is a space-time white noise on $[0,T]\times[0,L]$, with initial condition $u_{0}(x)=1$, Neumann, Dirichlet, or periodic boundary conditions. We assume that the coefficient $\sigma\in C^{1}(\mathbb{R})$ satisfies global Lipschitz and linear growth conditions.
Following [W86], this equation has a unique mild solution, which is adapted to the filtration generated by $W$, such that $E[u(t,x)^{2}]<\infty$ and it satisfies the evolution equation
@@
We assume the following technical condition, which is required for the proof of the central limit theorem.
Assumption 1.
There exists some positive valued measurable function $f_{\sigma}$ such that for any $T>0$, $f_{\sigma}\in L^{1}([0,T])$ and for all $t\in[0,T]$,
In this paper, we are interested in the asymptotic behavior as $L$ tends to infinity of the quantity
$F_{L}(t):=\frac{1}{L}\int_{0}^{L}\left\{u(t,x)-E[u(t,x)]\right\}\ \mathrm{d}x,$ (3)
where $L\geq 1$, $u(t,x)$ is the solution to (1).
Recall that the total variation distance between two random variables $X$ and $Y$ is defined as
$d_{\text{TV}}(X,Y)=\sup_{B\in\mathcal{B}(\mathbb{R})}\left|P(X\in B)-P(Y\in B)\right|$
where $\mathcal{B}(\mathbb{R})$ denotes the Borel $\sigma$-algebra with respect to the Euclidean topology on $\mathbb{R}$. We also write $d_{\text{TV}}(F,\mathcal{N}(0,1))$ for the total variation distance between the law of $F$ and the standard normal law.
@@
Suppose that $\operatorname{Var}(F_{L}(t))>0$. Then, for every $t>0$ there exists a real number $c=c(t)>0$ such that for all $L\geq 1$,
@@
We also prove the following functional central limit theorem.
@@
Fix $T>0$. Then, as $L\to\infty$,
@@
This kind of asymptotic behavior of spatial averages for stochastic partial differential equations has been studied extensively in recent years. This line of research was initiated by Huang, Nualart, and Viitasaari in [HNV20], where they established quantitative central limit theorem and its functional version for one-dimensional nonlinear stochastic heat equations driven by space-time white noise using Malliavin–Stein’s method (cf. [NP12]). After this result, similar central limit theorems for stochastic partial differential equations have been established under various settings. On the other hand, many of these studies deal with SPDEs whose solutions exhibit spatial stationarity. In our case, as well as in [P22], we consider SPDEs on the interval $[0,L]$, and therefore the solutions are not stationary in space, which requires technical conditions such as Assumption 1 for the proof of the central limit theorem. Moreover, the stochastic cable equation studied in this paper can be seen as a generalization of the stochastic heat equation considered in [P22]. For these reasons, our study can be regarded as a generalization of [P22].
@@
The nonlinear stochastic cable equation emerges within the framework of mathematical models of neurons, where it governs the propagation of electrical signals along the cylindrical structure of neurons. Neurons, the fundamental units of the nervous system, operate through an intricate interplay of chemical, biological, and electrical processes. In a standard mathematical idealization, neurons are represented as long, thin cylinders, analogous to electrical cables, enabling the application of cable theory. For such a cylindrical structure spanning from $x=0$ to $x=L$, we denote the electrical potential at position $x$ and time $t$ by $u(t,x)$, focusing on variations along the spatial coordinate $x$. This potential $u(t,x)$ is governed by a system of non-linear PDE’s, called the Hodgkin-Huxley equations, but in certain ranges of values of $u$, they are well-approximated by the cable equation:
@@
where $\beta/2>0$ represents the rate of diffusion within the neuron and $\alpha\in\mathbb{R}$ is the rate at which ions leak across the membrance [KX93]. The surface of the neuron is covered with synapses, through which it receives impulses of current. If the current arriving at $(t,x)$ is denoted by $F(t,x)$, the system satisfies the inhomogeneous PDE:
@@
Even when the system is at rest, occasional random impulses may occur, implying that $F$ generally contains a stochastic component. In particular, Walsh studied the case where $F$ is compound Poisson process or space-time white noise [W81]. In [W86], the nonlinear stochastic cable equation was studied for the case $F=\sigma(u)\dot{W}$.
@@ 2 Preliminaries
We write $\|Z\|_{k}$ instead of $E[|Z|^{k}]^{1/k}$, for every $Z\in L^{k}(\Omega)$.
@@ 2.1 Clark–Ocone formula
Let $\mathcal{H}=L^{2}([0,T]\times\mathbb{R})$. The Gaussian family $\{X(h)\}_{h\in\mathcal{H}}$ formed by the stochastic integrals
@@
defines an isonormal Gaussian process on the Hilbert space $\mathcal{H}$. In this framework we can develop the Malliavin calculus [N06]. We denote by $D$ the Malliavin derivative operator. Let $\{\mathcal{F}_{s}\}_{s\in[0,T]}$ denote the filtration generated by the space-time white noise $\dot{W}$. We recall the following Clark–Ocone formula:
$F=E[F]+\int_{[0,T]\times\mathbb{R}}E[D_{s,z}F\mid\mathcal{F}_{s}]\ W(\mathrm{d}s,\mathrm{d}z)\mathrm{~{}a.s.},$
valid for every $\mathcal{F}_{T}$measurable random variable $F$ in the Sobolev space $\mathbb{D}^{1,2}$ in the sense of Malliavin derivative. By using Jensen’s inequality for conditional expectations, the above Clark–Ocone formula yields the following Poincaré-type inequality:
$\left|\operatorname{Cov}[F,G]\right|\leq\int_{0}^{T}\int_{\mathbb{R}}\|D_{s,z}F\|_{2}\|D_{s,z}G\|_{2}\ \mathrm{d}z\mathrm{d}s$ (6)
for all $\mathcal{F}_{T}$measurable random variables $F,G\in\mathbb{D}^{1,2}$.
2.2 The Malliavin–Stein’s method
Let us now introduce some preliminary results of Malliavin–Stein’s method [NP12]. Stein’s method is a probabilistic technique to allow one to get some bounds for the distance between two probability measures. Combining Stein’s method and the Malliavin calculus, we can obtain useful estimate to prove quantitative central limit theorems.
@@ 2.2 The Malliavin–Stein’s method
Let $F=\delta(v)$ for some $v\in\operatorname{Dom}(\delta)$ which is the $L^{2}(\Omega)$-domain of the adjoint of the Malliavin derivative operator. Suppose that $F\in\mathbb{D}^{1,2}$ and $E[F^{2}]=1$. Then we have
@@
In order to establish the functional CLT in Theorem 2, we also need the following multivariate counterpart of Proposition 2.1.
@@
Fix $m\geq 2$. Let $F=(F_{1},\dots,F_{m})$ be a random vector such that for every $i=1,\dots,m$, $F_{i}=\delta(v_{i})$ for some $v_{i}\in\operatorname{Dom}(\delta)$ and $F_{i}\in\mathbb{D}^{1,2}$. Let $Z$ be an $m$-dimensional centered Gaussian random vector with covariance matrix $(C_{i,j})_{1\leq i,j\leq m}$. Then, for every $h\in C^{2}(\mathbb{R}^{m})$ that has bounded second partial derivatives,
@@
2.3 Moments and Malliavin derivative of $u(t,x)$
@@
Let $u(t,x)$ be the solution to (1). Then for all $k\geq 2$,
$\sup_{L\geq 1}\sup_{(t,x)\in[0,T]\times[0,L]}\|u_{n}(t,x)\|_{k}\leq c_{T,k}<\infty.$ (7)
@@
$\|D_{s,y}u(t,x)\|_{k}\leq\begin{cases}C_{T,k}1_{[0,L]}(y)p_{\beta(t-s)}(x-y)&\mathrm{(Neumann/Dirichlet~{}case)}\\ C_{T,k}1_{[0,L]}(y)G_{t-s}(x-y)&\mathrm{(periodic~{}case)}\end{cases}$
@@ 3 Asymptotic behavior of the covariance
The following result provides the asymptotic behavior of the covariance function of the renormalized sequence of processes $F_{L}(t)$ as $L$ tends to infinity.
@@
For every $t_{1},t_{2}>0$,
$\lim_{L\to\infty}\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]=\int_{0}^{t_{1}\wedge t_{2}}e^{-\alpha(t_{1}+t_{2}-2s)}f_{\sigma}(s)\ \mathrm{d}s,$
where the function $f_{\sigma}$ is defined in Assumption 1.
@@
$\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]$
@@
Then, by Assumption 1, we have
@@
$=\int_{0}^{t_{1}\wedge t_{2}}e^{-\alpha(t_{1}+t_{2}-2s)}f_{\sigma}(s)\ \mathrm{d}s.$

### v2

@@ Abstract
We study one-dimensional nonlinear stochastic cable equations driven by a multiplicative space-time white noise. Using the Malliavin–Stein method, we prove a central limit theorem for the spatial average of the solution. The convergence is established in the total variation distance with mild conditions. We also establish a functional central limit theorem with a technical assumption. Furthermore, we show that this assumption holds in a special case.
@@ 1 Introduction and main results
In this paper, we consider the nonlinear stochastic cable equation
@@
on $0,T]\times[0,L]$ for some constants $\alpha\in\mathbb{R}$ and $\beta>0$, where $T>0$ is fixed, $\dot{W}$ is a space-time white noise on $0,T]\times[0,L]$, with initial condition $u_{0}(x)=1$, Neumann, Dirichlet, or periodic boundary conditions. We assume the coefficient $\sigma$ is global Lipschitz.
According to Walsh [W86], the above equation admits a unique mild solution, which is adapted to the filtration generated by $W$ and satisfies the condition $E[u(t,x)^{2}]<\infty$. The mild solution satisfies the following equation
@@
We study the large $L$ asymptotics of the spatial average $F_{L}(t)$ of the solution, given by
$F_{L}(t):=\frac{1}{L}\int_{0}^{L}\left\{u(t,x)-E[u(t,x)]\right\}\ \mathrm{d}x.$ (3)
For two random variables $X$ and $Y$, the total variation distance is defined as
$d_{\text{TV}}(X,Y)=\sup_{B\in\mathcal{B}(\mathbb{R})}|P(X\in B)-P(Y\in B)|,$
where the supremum is taken over all sets $B$ in the Borel $\sigma$-algebra $\mathcal{B}(\mathbb{R})$. We denote by $d_{\text{TV}}(F,\mathcal{N}(0,1))$ the total variation distance between the law of a random variable $F$ and the standard normal distribution.
@@
Suppose that $\sigma(1)\neq 0$. Then, for every $t>0$ there exists a real number $c=c(t)>0$ such that for all $L\geq 1$,
@@
REMARK $1.$
Condition $\sigma(1)\neq 0$ guarantees that $\operatorname{Var}(F_{L}(t))>0$. This follows from (11), (13), (14), (35), (36) and (37).
We assume the following technical condition, which is required for the proof of the functional central limit theorem.
Assumption 1.
There exists some nonnegative valued measurable function $f_{\sigma}$ such that for any $T>0$, $f_{\sigma}\in L^{1}([0,T])$ and for all $t\in[0,T]$,
We state the following functional central limit theorem.
@@
Fix $T>0$. Suppose that Assumption 1 holds. Then, as $L\to\infty$,
@@
In recent years, considerable attention has been paid to the asymptotic behavior of spatial averages for solutions to stochastic partial differential equations. This area of inquiry was first explored by Huang, Nualart, and Viitasaari in [HNV20]. Utilizing the Malliavin–Stein method (see [NP12]), they proved a central limit theorem, along with its functional counterpart, for the one-dimensional nonlinear stochastic heat equation with space-time white noise. Following their seminal work, analogous central limit theorems for various types of stochastic partial differential equations have been established. For example, Huang, Nualart, Viitasaari, and Zheng studied the $d$-dimensional stochastic heat equation with colored noise [HNVZ20]. There has been much research on the stochastic heat equation under different settings; see, for example, [ANTV22, BY23, CKNP23, KNP21, KY22, NXZ22, NZ20, P22]. Similar results are also known for the stochastic wave equation; see, for instance, [DNZ20] for the one-dimensional case, [GNZ21] for the two-dimensional case, and [E24, E25] for the three and higher dimensions. For other types of SPDEs, Liu and Shen studied a central limit theorem for the solution to an SPDE related to a pseudo-differential operator that generates a stable-like process [LS23].
The previous study most relevant to our setting is that of Pu [P22]. Pu considered the one-dimensional stochastic heat equation with boundary conditions on a bounded interval $[0,L]$ and analyzed the solution using its Wiener chaos decomposition. In particular, we remark that Assumption 1 is necessary for the proof of the functional central limit theorem when considering SPDEs on a bounded domain, as in the present work and that of Pu [P22].
@@
The nonlinear stochastic cable equation arises in the mathematical modeling of neurons, where it describes the propagation of electrical signals along their cylindrical structure. Neurons, the fundamental components of the nervous system, operate through a sophisticated interplay of chemical, biological, and electrical phenomena. In a common mathematical simplification, a neuron is idealized as a long, thin cylinder, similar to an electrical cable. For such a cylinder on the interval $[0,L]$, we assume the electrical potential $u(t,x)$ depends only on position $x\in[0,L]$ and time $t$. While this potential $u(t,x)$ is accurately governed by the Hodgkin–Huxley equations, for specific ranges of $u$, these equations can be closely approximated by the cable equation:
@@
with $\beta/2>0$ representing the diffusion rate within the neuron and $\alpha\in\mathbb{R}$ being the rate of ion leakage through the membrane [KX93]. The neuron’s surface receives current impulses through synapses. If this incoming current at $(t,x)$ is represented by $F(t,x)$, the system obeys the inhomogeneous PDE:
@@
Even in a resting state, occasional random impulses may occur, which implies that $F$ generally contains a stochastic component. For instance, Walsh studied the case where $F$ is a compound Poisson process or a space-time white noise [W81]. The nonlinear stochastic cable equation with $F=\sigma(u)\dot{W}$ was analyzed in [W86].
@@ 2 Preliminaries
For any $Z\in L^{k}(\Omega)$, we define $\|Z\|_{k}:=E[|Z|^{k}]^{1/k}$.
Throughout this paper, we assume that $\sigma$ is globally Lipschitz continuous with constant $K_{\sigma}$. This implies the following linear growth condition:
$|\sigma(y)|\leq M_{\sigma}(1+|y|)$
where $M_{\sigma}=\max\{K_{\sigma},|\sigma(0)|\}$.
@@ 2.1 Clark–Ocone formula
Let $\mathcal{H}$ be the Hilbert space $L^{2}([0,T]\times\mathbb{R})$. The family of stochastic integrals
@@
constitutes an isonormal Gaussian process $\{X(h)\}_{h\in\mathcal{H}}$. In this context, we employ the tools of Malliavin calculus (see, e.g., [N06]), and let $D$ denote the associated Malliavin derivative operator. Let $\{\mathcal{F}_{s}\}_{s\in[0,T]}$ be the filtration generated by the space-time white noise $\dot{W}$. A key result is the Clark–Ocone formula, which provides the following representation for any $\mathcal{F}_{T}$-measurable random variable $F$ in the Sobolev space $\mathbb{D}^{1,2}$:
$F=E[F]+\int_{[0,T]\times\mathbb{R}}E[D_{s,z}F\mid\mathcal{F}_{s}]\ W(\mathrm{d}s,\mathrm{d}z)\quad\text{a.s.}$
A well-known consequence of this formula, obtained via Jensen’s inequality for conditional expectations, is the following Poincaré-type inequality:
$\left|\operatorname{Cov}[F,G]\right|\leq\int_{0}^{T}\int_{\mathbb{R}}\|D_{s,z}F\|_{2}\|D_{s,z}G\|_{2}\ \mathrm{d}z\mathrm{d}s,$ (6)
which holds for any pair of $\mathcal{F}_{T}$-measurable random variables $F,G\in\mathbb{D}^{1,2}$.
2.2 The Malliavin–Stein method
We now recall some key results from the Malliavin–Stein method. This method merges the Malliavin calculus with Stein’s method to provide quantitative bounds on the distance between probability measures, proving particularly effective for central limit theorems [NP12].
@@ 2.2 The Malliavin–Stein method
Let $F=\delta(v)$ for some $v\in\operatorname{Dom}(\delta)$, where $\operatorname{Dom}(\delta)$ is the $L^{2}(\Omega)$-domain of the adjoint of the Malliavin derivative operator. Suppose that $F\in\mathbb{D}^{1,2}$ and $E[F^{2}]=1$. Then, the following inequality holds:
@@
The proof of the functional CLT in Theorem 2 requires a multivariate version of Proposition 2.1, which we state below.
@@
Fix an integer $m\geq 2$. Let $F=(F_{1},\dots,F_{m})$ be a random vector where each component has the form $F_{i}=\delta(v_{i})$ for some $v_{i}\in\operatorname{Dom}(\delta)$ and $F_{i}\in\mathbb{D}^{1,2}$. Let $Z$ denote an $m$-dimensional centered Gaussian random vector with a given covariance matrix $(C_{i,j})_{1\leq i,j\leq m}$. Then, for any function $h\in C^{2}(\mathbb{R}^{m})$ with bounded second partial derivatives, we have
@@
2.3 Moments and Malliavin derivative of the solution
@@
Fix $T>0$. Then for all $k\geq 2$, there exists $c_{T,k}>0$ such that
$\sup_{L\geq 1}\sup_{(t,x)\in[0,T]\times[0,L]}\|u(t,x)\|_{k}\leq c_{T,k}<\infty.$ (7)
@@
$\|D_{s,y}u(t,x)\|_{k}\leq\begin{cases}C_{T,k}1_{[0,L]}(y)p_{\beta(t-s)}(x-y)&\mathrm{(Neumann/Dirichlet~case)}\\ C_{T,k}1_{[0,L]}(y)G_{t-s}(x,y)&\mathrm{(periodic~case)}\end{cases}$
@@ 3 Asymptotic behavior of the covariance
The following lemmas are prepared for the proof of Proposition 3.1.
Lemma 3.2.
In the case of Neumann or periodic boundary conditions,
$\lim_{t\to 0}\sup_{L\geq 1}\sup_{x\in[0,L]}|E[\sigma(u(t,x))^{2}]-\sigma(1)^{2}|=0.$
Proof.
Fix $T>0$. For every $(t,x)\in[0,T]\times[0,L]$,
$|E[\sigma(u(t,x))^{2}]-\sigma(1)^{2}|$
$=|E[(\sigma(u(t,x))-\sigma(1))(\sigma(u(t,x))+\sigma(1))]|$
$\leq\sqrt{E[|\sigma(u(t,x))-\sigma(1)|^{2}]}\sqrt{E[|\sigma(u(t,x))+\sigma(1)|^{2}]}$
$\leq\sqrt{E[K_{\sigma}^{2}|u(t,x)-1|^{2}]}\sqrt{E[2\sigma(u(t,x))^{2}+2\sigma(1)^{2}]}$
$\leq K_{\sigma}\sqrt{E[|u(t,x)-1|^{2}]}\sqrt{4M_{\sigma}^{2}(c_{T,2}^{2}+1)+2\sigma(1)^{2}},$
where $K_{\sigma}$ is the Lipschitz constant of $\sigma$, and $M_{\sigma}$ is the constant in the linear growth condition. From (2) and Itô’s isometry, we get
$E[|u(t,x)-1|^{2}]$ (8)
$\leq 2\left|\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y-1\right|^{2}+2\int_{0}^{t}\int_{0}^{L}G_{t-s}(x,y)^{2}E[\sigma^{2}(u(s,y))]\ \mathrm{d}s\mathrm{d}y$
$\leq 2\left|\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y-1\right|^{2}+4M_{\sigma}^{2}(1+c_{T,2}^{2})\int_{0}^{t}\int_{0}^{L}G_{t-s}(x,y)^{2}\ \mathrm{d}s\mathrm{d}y.$
Since we assume Neumann or periodic boundary conditions, using (38) and semigroup property of $G$, we have
$E[|u(t,x)-1|^{2}]$
$\leq 2|e^{-\alpha t}-1|^{2}+4M_{\sigma}^{2}(1+c_{T,2}^{2})\int_{0}^{t}G_{2(t-s)}(x,x)\ \mathrm{d}s$
In the case of Neumann boundary conditions, using (39), we get
$\int_{0}^{t}G_{2(t-s)}(x,x)\ \mathrm{d}s$ $\leq K_{2T}e^{2|\alpha|T}\int_{0}^{t}\frac{1}{\sqrt{4\pi\beta(t-s)}}\ \mathrm{d}s$
$=K_{2T}e^{2|\alpha|T}\sqrt{\frac{t}{\pi\beta}}.$
Then, we have
$E[|u(t,x)-1|^{2}]$
$\leq 2|e^{-\alpha t}-1|^{2}+4M_{\sigma}^{2}(1+c_{T,2}^{2})K_{2T}e^{2|\alpha|T}\sqrt{\frac{t}{\pi\beta}}$
In the case of periodic boundary conditions, using (37) and the following identity [P22]*A.17
$\sum_{j\in\mathbb{Z}}p_{r}(j)=\sum_{n\in\mathbb{Z}}e^{-2r\pi^{2}n^{2}}\quad\text{for}~\text{all}~r>0,$
we have
$\int_{0}^{t}G_{2(t-s)}(x,x)\ \mathrm{d}s$
$=\int_{0}^{t}e^{-2\alpha(t-s)}\sum_{j\in\mathbb{Z}}p_{2\beta(t-s)}(jL)\ \mathrm{d}s$
$\leq\int_{0}^{t}e^{-2\alpha(t-s)}\sum_{j\in\mathbb{Z}}p_{2\beta(t-s)}(j)\ \mathrm{d}s$
$=\int_{0}^{t}e^{-2\alpha(t-s)}\sum_{n\in\mathbb{Z}}e^{-4\beta(t-s)\pi^{2}n^{2}}\ \mathrm{d}s$
$\leq e^{2|\alpha|T}\sum_{n\in\mathbb{Z}}\int_{0}^{t}e^{-4\beta(t-s)\pi^{2}n^{2}}\ \mathrm{d}s$
$=e^{2|\alpha|T}\sum_{n\in\mathbb{Z}}\frac{1-e^{-4\beta t\pi^{2}n^{2}}}{4\beta\pi^{2}n^{2}}.$
Then,
$E[|u(t,x)-1|^{2}]$
$\leq 2|e^{-\alpha t}-1|^{2}+4M_{\sigma}^{2}(1+c_{T,2}^{2})e^{2|\alpha|T}\sum_{n\in\mathbb{Z}}\frac{1-e^{-4\beta t\pi^{2}n^{2}}}{4\beta\pi^{2}n^{2}}.$
Applying the Lebesgue dominated convergence theorem, we get
$\lim_{t\to 0}\sum_{n\in\mathbb{Z}}\frac{1-e^{-4\beta t\pi^{2}n^{2}}}{4\beta\pi^{2}n^{2}}=\sum_{n\in\mathbb{Z}}\lim_{t\to 0}\frac{1-e^{-4\beta t\pi^{2}n^{2}}}{4\beta\pi^{2}n^{2}}=0.$
Therefore, in the case of Neumann or periodic boundary conditions,
$\lim_{t\to 0}\sup_{L\geq 1}\sup_{x\in[0,L]}|E[\sigma(u(t,x))^{2}]-E[\sigma(1)^{2}]|=0.$
Lemma 3.3.
In the case of Dirichlet boundary conditions,
$\lim_{t\to 0}\sup_{L\geq 1}\left|\frac{1}{L}\int_{0}^{L}E[\sigma(u(t,x))^{2}]\ \mathrm{d}x-\sigma(1)^{2}\right|=0.$
Proof.
Fix $T>0$. For every $(t,x)\in[0,T]\times[0,L]$,
$\left|\frac{1}{L}\int_{0}^{L}E[\sigma(u(t,x))^{2}]\ \mathrm{d}x-\sigma(1)^{2}\right|$
$=\left|\frac{1}{L}\int_{0}^{L}E[(\sigma(u(t,x))-\sigma(1))(\sigma(u(t,x))+\sigma(1))]\ \mathrm{d}x\right|$
$\leq\sqrt{\frac{1}{L}\int_{0}^{L}E[|\sigma(u(t,x))-\sigma(1)|^{2}]\ \mathrm{d}x}\sqrt{\frac{1}{L}\int_{0}^{L}E[|\sigma(u(t,x))+\sigma(1)|^{2}]\ \mathrm{d}x}$
$\leq\sqrt{\frac{1}{L}\int_{0}^{L}E[K_{\sigma}^{2}|u(t,x)-1|^{2}]\ \mathrm{d}x}\sqrt{\frac{1}{L}\int_{0}^{L}E[2\sigma(u(t,x))^{2}+2\sigma(1)^{2}]\ \mathrm{d}x}$
$\leq K_{\sigma}\sqrt{\frac{1}{L}\int_{0}^{L}E[|u(t,x)-1|^{2}]\ \mathrm{d}x}\sqrt{4M_{\sigma}^{2}(c_{T,2}^{2}+1)+2\sigma(1)^{2}}.$
From (2) and Itô’s isometry, we get (8). Since we assume Dirichlet boundary conditions, using semigroup property of $G$ and (39), we have
$E[|u(t,x)-1|^{2}]$
$\leq 2\left|\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y-1\right|^{2}+4M_{\sigma}^{2}(1+c_{T,2}^{2})\int_{0}^{t}G_{2(t-s)}(x,x)\ \mathrm{d}s$
$\leq 2\left|\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y-1\right|^{2}+4M_{\sigma}^{2}(1+c_{T,2}^{2})K_{2T}e^{2|\alpha|T}\int_{0}^{t}\frac{1}{\sqrt{4\pi\beta(t-s)}}\ \mathrm{d}s$
$\leq 2\left|\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y-1\right|^{2}+4M_{\sigma}^{2}(1+c_{T,2}^{2})K_{2T}e^{2|\alpha|T}\sqrt{\frac{t}{\pi\beta}}.$
By (36), we get
$\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y$
$=2e^{-\alpha t}\sum_{n=1}^{\infty}\sin\left(\frac{n\pi x}{L}\right)\frac{1-(-1)^{n}}{n\pi}e^{-\frac{n^{2}\pi^{2}\beta t}{2L^{2}}}.$
Then, applying the $L^{2}([0,L])$-orthogonality of the functions $\{x\mapsto\sin(n\pi x/L)\}_{n=1}^{\infty}$,
$\frac{1}{L}\int_{0}^{L}\left|\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y-1\right|^{2}\ \mathrm{d}x$ (9)
$=4e^{-2\alpha t}\sum_{n=1}^{\infty}\frac{1}{L}\int_{0}^{L}\sin^{2}\left(\frac{n\pi x}{L}\right)\ \mathrm{d}x\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}e^{-\frac{n^{2}\pi^{2}\beta t}{L^{2}}}$
$\quad+1-4e^{-\alpha t}\sum_{n=1}^{\infty}\frac{1}{L}\int_{0}^{L}\sin\left(\frac{n\pi x}{L}\right)\ \mathrm{d}x\frac{1-(-1)^{n}}{n\pi}e^{-\frac{n^{2}\pi^{2}\beta t}{2L^{2}}}$
$=2e^{-2\alpha t}\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}e^{-\frac{n^{2}\pi^{2}\beta t}{L^{2}}}+1-4e^{-\alpha t}\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}e^{-\frac{n^{2}\pi^{2}\beta t}{2L^{2}}}.$
Now, for every $L\geq 1$,
$\left|\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}e^{-\frac{n^{2}\pi^{2}\beta t}{L^{2}}}-\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}\right|$
$=\left|\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}(e^{-\frac{n^{2}\pi^{2}\beta t}{L^{2}}}-1)\right|$
$\leq\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}|e^{-\frac{n^{2}\pi^{2}\beta t}{L^{2}}}-1|$
$\leq\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}|e^{-n^{2}\pi^{2}\beta t}-1|.$
Applying the dominated convergence theorem,
$\lim_{t\to 0}\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}|e^{-n^{2}\pi^{2}\beta t}-1|=0.$
Combining the above with the following identity
$\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}=\sum_{k=1}^{\infty}\frac{4}{(2k-1)^{2}\pi^{2}}=\frac{1}{2}$
yields
$\lim_{t\to 0}\sup_{L\geq 1}\left|\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}e^{-\frac{n^{2}\pi^{2}\beta t}{L^{2}}}-\frac{1}{2}\right|=0.$
Similarly, we get
$\lim_{t\to 0}\sup_{L\geq 1}\left|\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}e^{-\frac{n^{2}\pi^{2}\beta t}{2L^{2}}}-\frac{1}{2}\right|=0.$
These together with (9) imply
$\lim_{t\to 0}\sup_{L\geq 1}\frac{1}{L}\int_{0}^{L}\left|\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y-1\right|^{2}\ \mathrm{d}x=0.$
Therefore, we have
$\lim_{t\to 0}\sup_{L\geq 1}\frac{1}{L}\int_{0}^{L}E[|u(t,x)-1|^{2}]\ \mathrm{d}x=0.$
This completes the proof.
@@
The following results provide the asymptotic behavior of the covariance function of the renormalized sequence of processes $F_{L}(t)$ as $L$ tends to infinity.
@@
There exists $\delta>0$, for every $t_{1},t_{2}>0$,
$\liminf_{L\to\infty}\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]\geq\frac{\sigma(1)^{2}}{2}\int_{0}^{t_{1}\wedge t_{2}\wedge\delta}e^{-\alpha(t_{1}+t_{2}-2s)}\ \mathrm{d}s.$ (10)
@@
$\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]$ (11)
@@
Then, we have
$\liminf_{L\to\infty}\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]$ (12)
$=\liminf_{L\to\infty}\int_{0}^{t_{1}\wedge t_{2}}e^{-\alpha(t_{1}+t_{2}-2s)}\frac{1}{L}\int_{0}^{L}E[\sigma(u(s,z))^{2}]\ \mathrm{d}z\mathrm{d}s$
In the case of Neumann or periodic boundary conditions, from Lemma 3.2, there exists $\delta>0$, for every $L\geq 1$ and $(s,z)\in[0,\delta]\times[0,L]$,
$E[\sigma(u(s,z))^{2}]\geq\frac{\sigma(1)^{2}}{2}.$ (13)
In the case of Dirichlet boundary conditions, from Lemma 3.3, there exists $\delta>0$, for every $L\geq 1$ and $s\in[0,\delta]$,
$\frac{1}{L}\int_{0}^{L}E[\sigma(u(s,z))^{2}]\ \mathrm{d}z\geq\frac{\sigma(1)^{2}}{2}.$ (14)
These together with (12) implies (10). ∎
Proposition 3.2.
Suppose that Assumption 1 holds. For every $t_{1},t_{2}>0$,
$\lim_{L\to\infty}\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]=\int_{0}^{t_{1}\wedge t_{2}}e^{-\alpha(t_{1}+t_{2}-2s)}f_{\sigma}(s)\ \mathrm{d}s,$ (15)
where the function $f_{\sigma}$ is defined in Assumption 1.
Proof.
By the same argument as in the proof of Proposition 3.1, for every $t_{1},t_{2}>0$,
@@
This together with Assumption 1 implies (15). ∎